# A Rollout-Based Algorithm and Reward Function for Efficient Resource Allocation in Business Processes

### v1

@@ Abstract
Resource allocation plays a critical role in minimizing cycle time and improving the efficiency of business processes. Recently, Deep Reinforcement Learning (DRL) has emerged as a powerful tool to optimize resource allocation policies in business processes. In the DRL framework, an agent learns a policy through interaction with the environment, guided solely by reward signals that indicate the quality of its decisions. However, existing algorithms are not suitable for dynamic environments such as business processes. Furthermore, existing DRL-based methods rely on engineered reward functions that approximate the desired objective, but a misalignment between reward and objective can lead to undesired decisions or suboptimal policies. To address these issues, we propose a rollout-based DRL algorithm and a reward function to optimize the objective directly. Our algorithm iteratively improves the policy by evaluating execution trajectories following different actions. Our reward function directly decomposes the objective function of minimizing the mean cycle time. Maximizing our reward function guarantees that the objective function is minimized without requiring extensive reward engineering. The results show that our method consistently learns the optimal policy in all six evaluated business processes, outperforming the state-of-the-art algorithm that can only learn the optimal policy in two of the evaluated processes.
@@ 1 Introduction
Engineered rewards are designed to guide decision-making agents toward achieving a specific objective. These engineered rewards are approximations of the objective, and maximizing this reward should also optimize the objective. However, engineered rewards, when misspecified, can lead to unintended agent behavior [1]. One risk is reward hacking, where an agent exploits loopholes in the reward function, resulting in unintended or undesirable behaviors without completing the intended objective [15]. For example, a cleaning robot may be given points for picking up trash. Instead of cleaning up trash, it can also knock over trash to earn more points. This behavior maximizes the number of points but fails to achieve the true goal of cleaning an area. Similar unintended behaviors can arise in business processes where an agent optimizes a different objective than intended.
@@
In the domain of BPO, three different reward functions for cycle time minimization have been proposed, but each has inherent limitations that can lead to suboptimal policies. The first reward function rewards an agent +1 for each case completion but does not account for the realized cycle time of the case [9, 23]. The second reward function penalizes the agent equivalent to the cycle time of a completed case [5]. However, this approach may inadvertently incentivize the agent to delay or avoid completing cases to minimize penalties. The third reward function returns the inverse of the cycle time when a case completes, ensuring that fast case completions yield a higher positive reward [11, 12]. However, this reward function can be gamed by the agent, which can result in suboptimal policies. While these methods can lead to efficient policies for certain processes, they also require problem-specific engineering of reward functions and algorithms, leading to suboptimal policies or failure to learn a policy at all. Furthermore, the typical DRL algorithms, such as proximal policy optimization (PPO) [17], used in the aforementioned studies are not specifically designed to handle dynamic environments, such as a business process, making it challenging to learn effective policies [21].
@@
This paper introduces a rollout-based DRL algorithm and a reward function to optimize the mean cycle time. Our algorithm directly determines the sum of rewards following different actions, removing the need to approximate state-action values. Our reward function eliminates the need for extensive reward engineering, ensuring that the learned policies directly optimize the desired objective. We provide a comprehensive evaluation of our algorithm and benchmark it with existing reward functions using the state-of-the-art PPO algorithm [17]. The results show that our algorithm successfully learns the optimal policy across all six evaluated business processes. Our method outperforms the current state-of-the-art algorithm which only learns the optimal policy in two processes.
Against this background, the remainder of the paper is structured as follows. Section 2 provides an overview of the related work. Section 3 introduces concepts related to resource allocation in business processes and its optimization algorithms. Section 4 details our proposed learning algorithm and reward function. Furthermore, this section details how we can compute the optimal policy for each evaluation scenario. Section 5 shows the results of our method compared to the optimal policy. In this section, we also compare the performance of different reward functions. Last, Section 6 concludes our work.
@@ 2.1 Algorithms for business process optimization
Dynamic approaches assign resources at runtime using process state data. Park and Song [16] employ predictive modeling to minimize costs. Recently, RL has emerged as a powerful tool for BPO due to its ability to adapt to observed process dynamics. Early work applied Q-learning to minimize cycle time [5]. However, Q-learning struggles with scalability as it stores each state-action pair separately, making it inefficient for large-scale, continuous business environments. DRL deals with this issue with approximators [20] and has been used for optimizing business processes [23, 9, 10], demonstrating effectiveness in large-scale processes [11, 12].
Popular DRL algorithms have been successfully applied in game environments [14, 17]. However, these environments differ from BPO settings, as they often have well-defined, deterministic rules. For instance, in chess, a move has a predictable and deterministic outcome. In contrast, business processes are driven by random events, such as unpredictable case arrivals and variable activity durations. Standard DRL algorithms are not explicitly designed to handle such stochastic environments [21]. Within the BPO domain, applications of these algorithms often do not converge to optimal policies [5, 11, 12, 23]. To address this issue, we propose a rollout-based algorithm to learn optimal policies for business processes. Our approach is inspired by the algorithm of Lagoudakis and Parr [8] and extended to fit the requirements of resource allocation for business processes.
@@ 2.2 Reward functions for business process optimization
In BPO, three reward functions have been proposed for cycle time minimization: (1) a reward of +1 per completed case [9, 23]; (2) a penalty equal to cycle time per completed case [5]; and (3) a reward of $\frac{1}{1+CT}$ per completed case, where $CT$ is cycle time [11, 12]. While these reward functions encourage cycle time minimization, they are prone to reward hacking [1, 15], in which the agent optimizes a different objective than intended.
@@
In this paper, we introduce a reward function that decomposes the contribution to the sum of cycle times and integrate it with a rollout-based DRL algorithm to learn optimal resource allocation policies.
@@ 3 Background
This section introduces the resource allocation problem in business processes and details how it can be formulated as a sequential decision-making problem. Section 3.1 introduces key concepts related to resource allocation. Section 3.2 presents the Markov Decision Process (MDP) framework and how DRL can be applied to learn a resource allocation policy.
@@ 3.1 Resource allocation in business processes
Figure 1: Example of a process model with two activities, three resources, and four active cases. Activities A and B have unassigned activity instances ($K_{\textsf{A}}=\{3,4\},K_{\textsf{B}}=\{1\}$). The yellow resource is processing activity instance 2 (=$\mathcal{R}^{-}=\{r_{\texttt{yellow}}\}$, $B=\{(2,r_{\texttt{yellow}})\}$). The red and green resources are available (+=$\mathcal{R}^{+}=\{r_{\texttt{red}},r_{\texttt{green}}\}$).
@@
Using these components, we can form a resource-to-activity assignment $(r,act)$. Considering resource eligibility, we define the set of allowed assignments as $\mathcal{D}=\{(r,act)\mid act\in\mathcal{A},r\in\mathcal{R}_{act}\}$. We define the set of possible assignments at runtime as $D=\{(r,act)\mid(r,act)\in\mathcal{D},r\in\mathcal{R}^{+},|K_{act}|>0\}$. Making an assignment $(r,act)$ starts the execution of $(r,k)$, which is the assignment of resource $r\in R_{act}$ to an activity instance $k\in K_{act}$. The set of ongoing assignments is defined as $B\subseteq Z\times\mathcal{R}^{-}$, where $Z$ is the set of assigned activity instances.
@@
All information related to the current condition of the process can be represented as a business process execution state $(C,K,\mathcal{R}^{+},\mathcal{R}^{-},B,t)$, where $t\in T\subseteq[0,\infty)$ is a moment in time. $\mathcal{S}$ is the set of all possible business process execution states. In each execution state, the agent can decide either to make an assignment $(r,act)\in\mathcal{D}$ or to postpone the decision. After making a decision, the state transitions to the next state $(C^{\prime},K^{\prime},\mathcal{R}^{+\prime},\mathcal{R}^{-\prime},B^{\prime},t^{\prime})$ in two steps, as follows. In the first step, if the agent makes an assignment $(r,act)$, a resource $r$ is assigned to an unassigned activity instance $k\in K_{act}$, i.e., $(r,k)$. Following this, changes are made to the resource sets +=+$\mathcal{R}^{+\prime}=\mathcal{R}^{+}-\{r\}$ and =$\mathcal{R}^{-\prime}=\mathcal{R}^{-}\cup\{r\}$, the set of unassigned activity instances $K^{\prime}=K-\{k\}$, and the set of ongoing assignments $B^{\prime}=B\cup\{(r,k)\}$. If the agent chooses to postpone, nothing happens in this step. In the second step, an event is triggered. This can either be the arrival of a new case or the completion of an assigned activity instance $z\in Z$. We denote the set of events that is possible in a state after executing action $a\in D\cup\{\mathrm{postpone}\}$ as $E(s,a)$. These events all have a timing probability distribution associated with them, which models how long it will take before they occur (e.g., how long it will take from an action until the next case arrival). Since we assume exponential distributions, the timing distribution of event $e\in E(s,a)$ can be represented by a rate $\lambda_{e}$. When a new case $c^{\prime}$ arrives, this also generates an initial activity instance $k^{\prime}$, such that $C^{\prime}=C\cup c^{\prime}$ and $K^{\prime}=K\cup k^{\prime}$. When an assignment $(r,k)$ completes, the state changes such that $B^{\prime}=B-\{(r,k)\}$, +=+$\mathcal{R}^{+\prime}=\mathcal{R}^{+}\cup\{r\}$ and =$\mathcal{R}^{-\prime}=\mathcal{R}^{-}-\{r\}$. Subsequently, a new activity instance $k^{\prime}$ may be generated depending on the process model, which changes $K^{\prime}=K\cup k^{\prime}$. If the completed activity instance results in the completion of a case, this case is removed from the set of active cases $C^{\prime}=C-\{c\}$. The cycle time of a case $CT_{c}$ is equal to the difference between its arrival and departure time.
@@ 4.1 The business process as a CTMDP
  • A binary mapping $\mathcal{R}\rightarrow\{0,1\}$, indicating whether the resource is available or not. For resource $r\in\mathcal{R}$, this feature equals 1 if $r\in R^{+}$ and 0 if $r\in R^{-}$.
@@ 4.2 Rollout-based algorithm
The algorithm that we use to find the best policy for assigning resources to tasks is based on the rollouts of that policy. A rollout refers to the process of simulating an agent’s interaction with the environment (i.e., repeatedly performing an action and checking its effect) while following a specific policy for a specific number of decision steps. The benefit of using rollouts is that – if we define the reward function properly – we can use the sum of rewards, which is also called the return, to represent exactly the objective we aim to optimize, such as the sum of cycle times or the total resource allocation cost. We use a rollout to estimate the expected return of a policy, given by:
@@
where $R_{j}=R(s_{j},a_{j})$ is the reward of decision step $j$, and $N$ represents the number of decision steps in each rollout. To improve the reliability of the estimate, we perform $M$ rollouts and compute the average return. We use common random numbers to reduce variance across rollouts of different actions, ensuring each action is evaluated with the same sequence of random numbers [21]. Algorithm 1 presents our rollout-based approach.
@@
$\pi=\pi_{0}$;
$V^{\pi}(s_{0})\leftarrow\text{{evaluate}}(\mathcal{M},\pi)$
for $i=1,2,...,I$ do
       Generate states $\mathcal{S}$ from $\mathcal{M}$;
       dataset = $\emptyset$;
       foreach $s\in\mathcal{S}$ do
             foreach $a\in A(s)$ do
                   $G^{\pi}(s,a)\leftarrow\text{{rollout}}(\mathcal{M},\pi,M,N)$;
@@
             dataset $\leftarrow$ dataset $\cup\{(s,a^{*})\}$;
            $\pi^{\prime}\leftarrow\text{{learn}}(\pi,\text{dataset})$;
             $V^{\pi^{\prime}}(s_{0})\leftarrow\text{{evaluate}}(\mathcal{M},\pi^{\prime})$;
             if $V^{\pi^{\prime}}(s_{0})>V^{\pi}(s_{0})$ then
                   $\pi=\pi^{\prime}$;
Return $\pi$;
@@
Our rollout-based algorithm consists of three main steps: initialization, evaluation, and improvement. In the initialization step, we define a bootstrap policy $\pi_{0}$, for which we use a greedy policy. During the evaluation step, we deploy $\pi$ on the business process and evaluate its performance. In the improvement step, we explore alternative actions and update the policy to $\pi^{\prime}$. The evaluation and improvement steps are repeated for $I$ iterations.
@@
Figure 3: Rollouts for the example in Fig. 2. For each action $a\in A(s_{t})$, a rollout is performed. Consequently, the state transitions to state $s_{t+1}$. Next, a rollout of $N-1$ decision steps following policy $\pi$ is performed to estimate the return $G_{t}$ of each action.

### v2

# A Rollout-Based Algorithm and Reward Function for Resource Allocation in Business Processes

@@ Abstract
Resource allocation plays a critical role in minimizing cycle time and improving the efficiency of business processes. Recently, Deep Reinforcement Learning (DRL) has emerged as a powerful technique to optimize resource allocation policies in business processes. In the DRL framework, an agent learns a policy through interaction with the environment, guided solely by reward signals that indicate the quality of its decisions. However, existing algorithms are not suitable for dynamic environments such as business processes. Furthermore, existing DRL-based methods rely on engineered reward functions that approximate the desired objective, but a misalignment between reward and objective can lead to undesired decisions or suboptimal policies. To address these issues, we propose a rollout-based DRL algorithm and a reward function to optimize the objective directly. Our algorithm iteratively improves the policy by evaluating execution trajectories following different actions. Our reward function directly decomposes the objective function of minimizing the cycle time, such that trial-and-error reward engineering becomes unnecessary. We evaluated our method in six scenarios, for which the optimal policy can be computed, and on a set of increasingly complex, realistically sized process models. The results show that our algorithm can learn the optimal policy for the scenarios and outperform or match the best heuristics on the realistically sized business processes.
@@ 1 Introduction
Engineered rewards are designed to guide decision-making agents toward achieving a specific objective. These engineered rewards are approximations of the objective, and maximizing this reward should also optimize the objective. However, engineered rewards, when misspecified, can lead to unintended agent behavior [1]. One risk is reward hacking, where an agent exploits loopholes in the reward function, resulting in unintended or undesirable behaviors without completing the intended objective [16]. For example, a cleaning robot may be given points for picking up trash. Instead of cleaning up trash, it can also knock over trash, thus maximizing the reward it earns, but failing to achieve the objective of cleaning an area. Similar unintended behaviors can arise in business processes where an agent optimizes a different objective than intended.
@@
In the domain of BPO, three different reward functions for cycle time minimization have been proposed, but each has inherent limitations that can lead to suboptimal policies. The first reward function rewards an agent +1 for each case completion but does not account for the realized cycle time of the case [9, 25]. The second reward function penalizes the agent equivalent to the cycle time of a completed case [5]. However, this approach may inadvertently incentivize the agent to delay or avoid completing cases to minimize penalties. The third reward function returns the inverse of the cycle time when a case completes, ensuring that fast case completions yield a higher positive reward [11, 12]. However, this reward function can be gamed by the agent, resulting in suboptimal policies. While these methods can lead to efficient policies for certain processes, they also require problem-specific engineering of reward functions and algorithms, leading to suboptimal policies or failure to learn a policy. Furthermore, the typical DRL algorithms, such as proximal policy optimization (PPO) [19], used in the aforementioned studies are not specifically designed to handle dynamic environments, such as a business process, making it challenging to learn effective policies [23].
@@
This paper introduces a rollout-based DRL algorithm and a reward function to optimize the mean cycle time. Our algorithm directly determines the sum of rewards following different actions, removing the need to approximate state-action values. Our reward function eliminates the need for extensive reward engineering, ensuring that the learned policies directly optimize the desired objective. We provide a comprehensive evaluation of our algorithm and benchmark it with existing reward functions using the state-of-the-art PPO algorithm [19]. The results show that our algorithm successfully learns the optimal policy across for typical business process pattenrs and outperforms the PPO algorithm.
The remainder of the paper is structured as follows. Section 2 provides an overview of the related work. Section 3 introduces background concepts. Section 4 explains our algorithm and reward function. Section 5 evaluates our method and Section 6 concludes our work.
@@ 2.1 Algorithms for business process optimization
Dynamic approaches assign resources at runtime using process state data. Park and Song [17] employ predictive modeling to minimize costs. Recently, RL has emerged as a powerful tool for BPO due to its ability to adapt to observed process dynamics. Early work applied Q-learning to minimize cycle time [5]. However, Q-learning struggles with scalability as it stores each state-action pair separately, making it inefficient for large-scale, continuous business environments. DRL deals with this issue with approximators [22] and has been used for optimizing business processes [9, 10, 25], demonstrating effectiveness in large-scale processes [11, 12, 15].
Popular DRL algorithms have been successfully applied in game environments [14, 19]. However, these environments differ from BPO settings, as they often have well-defined, deterministic rules. For instance, in chess, a move has a predictable and deterministic outcome. In contrast, business processes are driven by random events, such as unpredictable case arrivals and variable activity durations. Standard DRL algorithms are not explicitly designed to handle such stochastic environments [23]. Within the BPO domain, applications of these algorithms often do not converge to optimal policies [5, 11, 12, 25]. To address this issue, we propose a rollout-based algorithm to learn optimal policies for business processes. Our approach is inspired by other rollout-based algorithms [8, 23] and we extended this idea to fit the requirements of resource allocation for business processes. For a more extensive review of the resource allocation algorithms, we refer to Pufahl et al. [18].
@@ 2.2 Reward functions for business process optimization
In BPO, three reward functions have been proposed for cycle time minimization: (1) a reward of +1 per completed case [9, 25]; (2) a penalty equal to cycle time per completed case [5, 15]; and (3) a reward of $\frac{1}{1+CT}$ per completed case, where $CT$ is cycle time [11, 12]. While these reward functions encourage cycle time minimization, they are prone to reward hacking [1, 16], in which the agent optimizes a different objective than intended.
@@
In this paper, we introduce a reward function that decomposes the contribution to the sum of cycle times and integrates it with a rollout-based DRL algorithm to learn optimal resource allocation policies.
@@ 3 Background
This section introduces the resource allocation problem and how it can be formulated and solved as an MDP. Section 3.1 defines resource allocation concepts. Section 3.2 presents the Markov Decision Process (MDP) framework and how DRL can be applied to learn a resource allocation policy on an MDP.
@@ 3.1 Resource allocation in business processes
Figure 1: Example of a process model with two activities, three resources, and four active cases. Activities A and B have unassigned activity instances ($K_{\textsf{A}}=\{3,4\},K_{\textsf{B}}=\{1\}$). The yellow resource is processing activity instance 2 ($\mathcal{R}^{-}=\{r_{\texttt{yellow}}\}$, $B=\{(2,r_{\texttt{yellow}})\}$). The red and green resources are available ($\mathcal{R}^{+}=\{r_{\texttt{red}},r_{\texttt{green}}\}$).
@@
Using these components, we can form a resource-to-activity assignment $(r,act)$. Considering resource eligibility, we define the set of allowed assignments as $\mathcal{D}=\{(r,act)\mid act\in\mathcal{A},r\in\mathcal{R}_{act}\}$. We define the set of possible assignments at runtime as $D=\{(r,act)\mid(r,act)\in\mathcal{D},r\in\mathcal{R}^{+},|K_{act}|>0\}$. Making an assignment $(r,act)$ starts the execution of $(r,k)$, which is the assignment of resource $r\in R_{act}$ to an activity instance $k\in K_{act}$. Then, $r$ and $k$ are taken from $\mathcal{R}^{+}$ and the set of assigned activity instances $K$, and instead put into $\mathcal{R}^{-}$ and set of assigned activity instances $Z$, and the set of ongoing assignments $B\subseteq Z\times\mathcal{R}^{-}$.
@@
The current condition of the process can be represented as an execution state $(C,K,\mathcal{R}^{+},\mathcal{R}^{-},B,t)$, where $t\in T\subseteq[0,\infty)$ is a moment in time. $\mathcal{S}$ is the set of all possible execution states. In each state, the agent can decide either to make an assignment $(r,act)\in\mathcal{D}$ or to postpone the decision. After making a decision, the state transitions to the next state $(C^{\prime},K^{\prime},\mathcal{R}^{+\prime},\mathcal{R}^{-\prime},B^{\prime},t^{\prime})$ in two steps, as follows. In the first step, if the agent makes an assignment $(r,act)$, a resource $r$ is assigned to an unassigned activity instance $k\in K_{act}$, i.e., $(r,k)$. Next, changes are made to the resource sets $\mathcal{R}^{+\prime}=\mathcal{R}^{+}-\{r\}$ and $\mathcal{R}^{-\prime}=\mathcal{R}^{-}\cup\{r\}$, the set of unassigned activity instances $K^{\prime}=K-\{k\}$, and the set of ongoing assignments $B^{\prime}=B\cup\{(r,k)\}$. If the agent chooses to postpone, nothing happens in this step. In the second step, an event is triggered. This can either be the arrival of a new case or the completion of an assigned activity instance $k\in Z$. We denote the set of events that is possible in a state $s$ after executing action $a\in D\cup\{\mathrm{postpone}\}$ as $E(s,a)$. These events all have a timing probability distribution associated with them, which models how long it will take before they occur (e.g., how long it will take from an action until the next case arrival). Since we assume exponential distributions, the timing distribution of event $e\in E(s,a)$ can be represented by a rate $\lambda_{e}$. When a new case $c^{\prime}$ arrives, this also generates an initial activity instance $k^{\prime}$, such that $C^{\prime}=C\cup\{c^{\prime}\}$ and $K^{\prime}=K\cup\{k^{\prime}\}$. When an assignment $(r,k)$ completes, the state changes such that $B^{\prime}=B-\{(r,k)\}$, $\mathcal{R}^{+\prime}=\mathcal{R}^{+}\cup\{r\}$ and $\mathcal{R}^{-\prime}=\mathcal{R}^{-}-\{r\}$. Subsequently, a new activity instance $k^{\prime}$ may be generated depending on the process model, which changes $K^{\prime}=K\cup k^{\prime}$. If the completed activity instance results in the completion of a case, this case is removed from the set of active cases $C^{\prime}=C-\{c\}$. The cycle time of a case $CT_{c}$ is equal to the difference between its arrival and departure time.
@@ 4.1 The business process as a CTMDP
  • A binary mapping $\mathcal{R}\rightarrow\{0,1\}$, indicating whether the resource is available or not. For resource $r\in\mathcal{R}$, this feature equals 1 if $r\in\mathcal{R}^{+}$ and 0 if $r\in\mathcal{R}^{-}$.
@@ 4.2 Rollout-based algorithm
The algorithm that we use to find the best policy for assigning resources to tasks is based on the rollouts of that policy. A rollout refers to the process of simulating an agent’s interaction with the environment (i.e., repeatedly performing an action and checking its effect) while following a specific policy for a specific number of decision steps. The benefit of using rollouts is that we can use the sum of rewards, which is also called the return, to represent exactly the objective we aim to optimize, such as the sum of cycle times. We use a rollout to estimate the expected return of a policy, given by:
@@
where $R_{j}=R(s_{j},a_{j})$ is the reward of decision step $j$, and $N$ represents the number of decision steps in each rollout. To improve the reliability of the estimate, we perform $M$ rollouts and compute the average return. We use common random numbers to reduce variance across rollouts actions [23]. Algorithm 1 presents our rollout-based approach.
@@
2 $V^{\pi}(s_{0})\leftarrow\text{{evaluate}}(\mathcal{M},\pi)$;
3for $i=1,2,\dots,I$ do
 4 Generate states $\mathcal{S}$ from $\mathcal{M}$;
 5 dataset = $\emptyset$;
 6 foreach $s\in\mathcal{S}$ do
    7 foreach $a\in A(s)$ do
       8 $G^{\pi}(s,a)\leftarrow\text{{rollout}}(\mathcal{M},\pi,M,N)$;
[…]
    10 dataset $\leftarrow$ dataset $\cup\{(s,a^{*})\}$;
[…]
 12 $V^{\pi^{\prime}}(s_{0})\leftarrow\text{{evaluate}}(\mathcal{M},\pi^{\prime})$;
 13 if $V^{\pi^{\prime}}(s_{0})>V^{\pi}(s_{0})$ then
    14 $\pi=\pi^{\prime}$;
15Return $\pi$;
@@ 14 $\pi=\pi^{\prime}$;
Our rollout-based algorithm has three steps: initialization, evaluation, and improvement. In the initialization step (line 1), we define a bootstrap policy $\pi_{0}$, for which we use a greedy policy. During the evaluation step (lines 2, 12-14), we deploy $\pi$ on the business process and evaluate its performance. In the improvement step (lines 3-11), we explore alternative actions and update the policy to $\pi^{\prime}$. The evaluation and improvement steps are repeated for $I$ iterations.
@@
Figure 3: Rollouts for the example in Fig. 2, representing lines 7-10 of Algorithm 1. For each action $a\in A(s_{t})$, a rollout is performed. Consequently, the state transitions to state $s_{t+1}$. Next, a rollout of N1N-1 decision steps following policy $\pi$ is performed to estimate the return $G_{t}$ of each action.